# JailbreakLens: Interpreting Jailbreak Mechanism in the Lens of Representation and Circuit

### v1

@@ 4.3 Interpretation Aggregation
     $y_{i}\leftarrow\mathcal{M}(x+Y)$triangleright generate the next token
     $Y\leftarrow[Y,y_{i}]$triangleright add a generated token to the sequence
     $i\leftarrow i+1$triangleright update the token index
end while
$E_{R}\leftarrow\{E_{R}^{(1)},E_{R}^{(2)},...,E_{R}^{(i)}\}$
$E_{C}\leftarrow\{(rs_{S_{+}}^{(1)},rs_{S_{-}}^{(1)}),(rs_{S_{+}}^{(2)},rs_{S_{-}}^{(2)}),...,(rs_{S_{+}}^{(i)},rs_{S_{-}}^{(i)})\}$
Return: $E_{R},E_{C}$
5 Experiments
@@
The experiments are organized as follows. Firstly we introduce datasets and models used in this paper in Sec. 5.1. , as well as the detailed implements in Sec. 5.2. Next, we present the experimental results on how jailbreak prompts affect model’s internal representations in Sec. 5.3. Then we demonstrate and analyze the results on the impact of jailbreak prompts on model components including attention heads and MLP layers in Sec. 5.4. Finally, we present the results of examining the dynamic evolution of internal representations and key components, i.e., aggregated interpretation, during the token generation process in response to jailbreak prompts in Sec. 5.5.
5.1 Models and Datasets
@@
We conduct our experiment with four mainstream LLMs, i.e., Llama-2-7b-chat-hf, Llama-2-13b-chat-hf, Vicuna-7b-v1.5, and Vicuna-13b-v1.5.
@@
We construct a paired dataset consisting of harmful prompts and safe prompts. We choose Advbench[6] as the harmful dataset. For safe prompts, we construct a paired dataset where each sample pair contains a harmful prompt (e.g., how to steal personal information) from Advbench and a safe prompt (e.g., how to protect personal information) semantically identical to the harmful sample modified with GPT-4. From the above two datasets, we randomly select 400 samples, setting the test size to 0.3 for probe training.
@@
We analyze seven famous jailbreak methods for the experiment, with a brief introduction in Tab. I, including the method category in JailbreakZoo [11] and attack success rate on Llama2-7b. The typical example of each jailbreak strategy is shown in Appendix. A. The success of jailbreak evaluated with GenerativeJudge [40] with human check again. Moreover, the attack success rate (ASR) of these jailbreak strategies on all models we used in this paper is summarized in Tab. V in Appendix. A.
Jailbreak Category ASR on Llama2-7b
ReNellm[13] Rule 34%
CodeChaeleom[9] Rule 32%
DeepInception[8] Demonstration 32%
PAIR[10] Multi-agent 21%
GPTFuzz[7] Evolution 16%
GCG[6] Gradient 4%
Autodan[12] Evolution 4%
TABLE I: Jailbreak methods analyzed in the paper, with their categories and attack success rate on Llama2-7b.
5.2 Implementations
For representation analysis, we utilize three types of probes, i.e., linear-based, cluster-based and PCA-based, to analyze whether the model can effectively recognize the harmfulness of jailbreak prompts.
@@ 5.2 Implementations
We obtain the hidden states of the prompts in the dataset from each model layer and use the last token of each prompt as the input for the probes, the last token is particularly informative as it captures the accumulated context from all preceding tokens.
5.3 Results of Representation Analysis
In this section, we present the results of the analysis as well as the findings on the jailbreak mechanism at the representation level. Specifically, we conduct experiments from the perspectives including ➀ LLM Harmfulness Perception, ➁ Detecting Harmfulness in Jailbreak Prompts, ➂ Decoding Jailbreak Representations, and ➃ Impact of Model Scale on Jailbreak Vulnerability.
To present ➀ LLM harmfulness perception, we investigate the extent to which safety concepts are encoded internally in the model by training different types of probing classifiers to evaluate how well representations of safe prompts are distinguished from representations of harmful prompts. Then to present ➁ detecting harmfulness in jailbreak prompts, we classify jailbreak representations with well-trained probes to determine whether the model can identify the harmfulness within jailbreak representations. Moreover, for ➂ decoding jailbreak representations, we convert each model layer’s representation of jailbreak prompts into human-understandable words with logitlen to observe how jailbreak prompts pass through each model layer to achieve bypass safeguard. Finally, for ➃ impact of model scale, we compare the performance of jailbreak prompts in model representation space for models of different sizes to analyze the impact of model size on jailbreak.
5.3.1 LLM Harmfulness Perception
(a) Prediction accuracy of probes in each model. The accuracy of all types of probes approaches 99% after the early few layers.
(b) Probing results on different jailbreak strategies on each model. Most jailbreak representations are predicted as safe by the probes in the early layers, indicating all methods deceive the model to treat jailbreak prompt containing harmful intent as safe ones. while their deception varies in the deep layers.
@@ 5.3.1 LLM Harmfulness Perception
Figure 2: Prediction accuracy of the probes and probing results of different jailbreak methods.
@@
The prediction accuracy on the test dataset of three probes on LLMs is shown in Fig. 2(a). Taking the Llama2-7b model as an example, after the middle layers (around the $10_{th}$ layer), the prediction accuracy of all probes approaches 99%, and the linear probe slightly outperforms the other two. These results demonstrate that the model encodes safety concepts within its hidden states, enabling it to effectively distinguish between safe and harmful prompts.
To further validate whether the directions identified by probes represent safety, we project these direction vectors into the vocabulary space and examine the top probability words corresponding to both the positive and negative direction vectors. The results for the three types of probes trained on Llama2-7b are summarized in Fig. II.
@@
As shown in Tab. II, few decoded words of the linear probe’s contain affirmation (e.g., ‘Yes’ for positive direction) or refusal (e.g., ‘not’ for negative direction) meaning with other words lack clear interpretability. However, the PCA and cluster probes provide clearer results, i.e., the positive direction consistently decodes into words that imply affirmation (e.g., ‘Sure’ and ‘certain’), while the negative direction produces words that express refusal (e.g., ‘cannot’ and ‘Sorry’). This strongly validates that the directions identified by these probes correspond to safety-related concepts. Thus, the results further demonstrate that the model’s internal representations contain an inherent sense of safety, which can be effectively captured by the probes.
@@
TABLE II: Decoding the direction vectors captured by the probes in the vocabulary space. The positive direction consistently decodes into words that imply affirmation, while the negative direction produces words that express refusal.
Observation 1. The probing results reveal that the aligned model’s representations contain a discernible understanding of safety, with specific directions corresponding to affirmation or refusal, indicating an embedded capacity to differentiate between safe and harmful prompts.
5.3.2 Detecting Harmfulness in Jailbreak Prompts
@@
The results of probing jailbreak representations of each layer are shown in Fig. 2(b). As shown in Fig. 2(b), most jailbreak methods maintain a high probability of being predicted as safe by the probes in the few early layers (i.e., around the first 10 layers). This suggests that although the jailbreak strategies vary, their core mechanism lies in deceiving the model’s perception of harmfulness, i.e., making the model treat jailbreak prompts containing malicious intent as safe.
@@
However, different jailbreak methods vary in their ability to deceive the model’s harmfulness detection in the deep layers (i.e., around the last 10 layers). For example, in the case of Llama2-7b, as shown in the first row in Fig. 2(b), rule-based and demonstration-based methods such as ReNellm, CodeChameleon, and DeepInception are more deceptive than other methods, as these prompts are identified as safe across all the layers. In contrast, evolution-based methods and multi-agent methods like Autodan and PAIR are less deceptive. Since these methods initially deceive the model into perceiving their prompts as safe in the early layers but still recognized as harmful by the final layers.
@@
Observation 2. Jailbreak prompts bypass safeguard by deceiving model’s harmfulness perception but varying in effectiveness. Rule-based and demonstration-based methods are more deceptive, maintaining safe predictions even in the final layers, while other methods gradually identify their harmful intent as the model processes deeper layers.
5.3.3 Decoding the Jailbreak Representations
We first investigate how Llama2-7b generates tones when processing normal safe and harmful prompts by projecting the corresponding representations from each model layer into the vocabulary space. Since most words decoded from early layer representations are meaningless, we only display and analyze the decoding results of later layers, illustrated in Fig. 3. As shown in Fig. 3(a), the model’s representation of safe prompts in the middle layer (around the $18_{th}$ layer) consistently maps to affirmative words such as ‘Sure’ with high probability, without any refusal words like ‘Sorry’ appearing throughout the process. For harmful prompts, as demonstrated in Fig. 3(b), the trend is reversed, with the decoded words consistently carrying meanings of refusal.
(a) Safe prompt.
(b) Harmful prompt.
(c) DeepInception prompt.
(d) ReNellm prompt.
(e) GCG prompt.
(f) PAIR prompt.
@@ 5.3.3 Decoding the Jailbreak Representations
Figure 3: Decoding the jailbreak representation of each model layer in Llama2-7b in the vocabulary space.
@@
Then we explore how the tone generation for jailbreak prompt responses differs from normal prompts. For rule-based and demonstration-based methods, such as DeepInception and ReNellm (with Fig. 3(c) and Fig. 3(d) as examples), the initial word predictions for most prompts tend to produce words that align with the instruction, including task-specific terms like ‘def’ and ‘Layer’, as well as affirmative words like ‘Here’ and ‘Sure’. Notably, no refusal words appear throughout the decoding process. The decoding behavior of these jailbreak prompts closely resembles that of safe prompts, indicating that these prompts successfully bypass model’s harmfulness perception at representation level.
@@
For some prompts using gradient-based methods, the generation tone is not consistent. Fig. 3(e) shows an example of a GCG prompt whose decoding results match this pattern, where affirmation words such as ‘Sure’ and refusal words such as ‘Sorry’ both appear simultaneously in the top-5 predictions, although affirmation words dominate with significantly higher probabilities in the last few layers. This suggests that while these jailbreak prompts do not fully deceive the model, i.e., refusal words still appear with high probability, model’s representations are sufficiently perturbed, with affirmation words outweighing refusal words in the final layers.
@@
In contrast, for a few rule-based and multi-agent-based jailbreak prompts, with an example shown in Fig. 3(f), no affirmation words are decoded from the hidden states. Refusal words maintain a high probability from the middle layers to the final layer, closely resembling the behavior of harmful prompts. This indicates that these jailbreak prompts fail to deceive the model at the first token generation, i.e., the model recognizes the harmful nature of the prompts.
@@
Observation 3. The representations of highly deceptive prompts maintain a consistent tone, with the decoded tokens that have the highest probability all being affirmative. In contrast, less deceptive prompts do not exhibit this consistency, as the decoded tokens include both affirmative and negative tokens.
5.3.4 Impact of Model Scale on jailbreak Vulnerability
Furthermore, we analyze the impact of model scale on jailbreak mechanism at the representation level. By examining results from Llama2 (the $1_{st}$ and $2_{nd}$ rows in Fig. 2(b)) and Vicuna1.5 (the $3_{rd}$ and $4_{th}$ rows in Fig. 2(b)) series across different model scales (e.g., 7b vs. 13b), we observe that jailbreak methods maintain similar patterns in bypassing safety probes across models with different scales.
This persistent vulnerability indicates that while scaling up model parameters enhances representational power, it does not inherently strengthen the model’s resilience against attacks exploiting alignment weaknesses. Specifically, the underlying vulnerabilities targeted by these attacks remain deeply embedded in models of all sizes, as evidenced by that jailbreak success rate does not decrease obviously in models with larger scale. Therefore, parameter scaling alone offers limited defense against these jailbreak attacks, underscoring the necessity for alternative strategies, such as advanced training protocols or enhanced security mechanisms, to bolster model safety beyond mere size increases.
Observation 4. The alignment generalization capacity of the model does not show substantial improvement with increased parameter scale.
5.4 Results of Circuit Analysis
In this section, we present the experimental results and findings on the jailbreak mechanism at the circuit level. Specifically, we conduct experiments from the following perspectives: ➀ Key Components Location, ➁ Effect of Jailbreak Prompts on Key components, and ➂ Impact of Model Fine-tuning on Key Components Location.
For ➀ Key Components Location, we firstly analyze the behavior of model components on harmful and safe prompts and identify the components that have significant contribution to model safeguard via logit difference attribution introduced in Alg. 1. Then we convert the activation of key components into vocabulary space to validate whether the functionality of the identified components is associated with generation safety. For ➁ Effect of Jailbreak Prompts on Key Components, we investigate how these key components are impacted by the jailbreak prompts. Finally, for ➂ Impact of Model Fine-tuning, we compare the behaviors of key components on jailbreak prompts in Llama-2 models and Vicuna-1.5 models, which is obtained by instruction tuning from Llama-2 models. For visual clarity, only the results on the Llama2-7b and Vicuna1.5-7b models are presented in this section. More detailed experimental results on other models are shown in Appendix. B.1.
5.4.1 Key Components Location
To identify the main components that contribute to generating refusal response, Then we visualize the direct logit attribution defined in Eq. (2) of each attention head and MLP layer in Llama2-7b and Vicuna1.5-7b, where the results are shown in Fig. 4 and Fig. 5(a), respectively.
@@ 5.4.1 Key Components Location
Figure 4: Average refusal score for each attention head in Llama2-7b and Vicuna1.5-7b when responding to harmful and safe prompts. In both models, L21H14 contributes the most to refusals on harmful prompts, while L26H04 shows the highest contribution to affirmations on safe prompts.

### v2

@@ 3       $E_{R}^{(i)}\leftarrow\mathcal{P}(A_{l}^{(i)})$
       $y_{i}\leftarrow\mathcal{M}(x+Y)$ // generate the next token
       $Y\leftarrow[Y,y_{i}]$ // add the new token to the sequence
       $i\leftarrow i+1$ // update the token index
6
7 end while
8
$E_{R}\leftarrow\{E_{R}^{(1)},E_{R}^{(2)},...,E_{R}^{(i)}\}$ $E_{C}\leftarrow\{(rs_{S_{+}}^{(1)},rs_{S_{-}}^{(1)}),(rs_{S_{+}}^{(2)},rs_{S_{-}}^{(2)}),...,(rs_{S_{+}}^{(i)},rs_{S_{-}}^{(i)})\}$ return $E_{R},E_{C}$
Algorithm 2 Tracking the Evolution of Representation and Circuit
5. Experiments
@@
The experiments are organized as follows. We first introduce the experimental settings, including datasets, evaluated models, jailbreak methods and probes in Sec. 5.1. Then we present the experimental results on how jailbreak prompts affect model’s internal representations in Sec. 5.2. Then we demonstrate and analyze the results on the impact of jailbreak prompts on model components including attention heads and MLP layers in Sec. 5.3. Finally, we present the results on correlation analysis between representation and circuit performance and then provide joint and dynamic interpretation results, i.e., aggregated interpretation, during the token generation process in response to jailbreak prompts, in Sec. 5.4.
5.1. Experimental Settings
@@
We conduct our experiment with five mainstream LLMs, i.e., Llama-2-7b-chat-hf, Llama-2-13b-chat-hf, Llama-3-8b-Instruct, Vicuna-7b-v1.5, and Vicuna-13b-v1.5.
@@
For evaluation dataset, we construct a paired dataset consisting of harmful prompts from Advbench(Zou et al., 2023) and corresponding safe prompts, where each safe prompt was semantically modified from its harmful counterpart using GPT-4 (e.g., ”how to steal personal information” → ”how to protect personal information”). We randomly select 400 samples and set the test size to 0.3 for probe training.
@@
For jailbreak methods, we analyze seven mainstream jailbreak strategies for the experiment. The typical example of each jailbreak strategy is shown in Appendix. A.2. Successful jailbreak prompts identified by GenerativeJudge (Li et al., 2023a) are manually verified. The attack success rate (ASR) of these jailbreak strategies on all tested models are summarized in Appendix Table 4.
@@
For representation analysis, we utilize three types of probes, i.e., linear-based, cluster-based and PCA-based, to analyze whether the model can effectively recognize the harmfulness of jailbreak prompts, The detailed description of each probe technique is shown in Appendix. A.1 . We obtain the hidden states of the prompts in the dataset from each model layer and use the last token of each prompt as the input for probing, the last token is particularly informative as it captures the accumulated context from all preceding tokens.
(a) Prediction accuracy of probes in each model.
(b) Probing results of different jailbreak strategies.
@@
Figure 2. Prediction accuracy of the probes and probing results of different jailbreak methods.
5.2. Results of Representation Analysis
In this section, we present the experimental results and the findings on the jailbreak mechanism at the representation level. Specifically, we conduct experiments from the perspectives including ➀ LLM Harmfulness Perception, ➁ Probing Deceptive Harmfulness in Jailbreak Prompts, ➂ Decoding Jailbreak Representations.
To present ➀ LLM harmfulness perception, we investigate the degree to which safety concepts are encoded internally in the model by training different types of probing classifiers to evaluate how well representations of safe prompts are distinguished from representations of harmful prompts. For ➁ probing deceptive harmfulness in jailbreak prompts, we apply these trained probes to jailbreak representations to determine whether the model can identify the harmfulness within jailbreak representations. Moreover, for ➂ decoding jailbreak representations, we convert representation from each model layer of jailbreak prompts into human-interpretable words with Logitlen to observe how jailbreak prompts progress through layers to bypass safeguards.
5.2.1. LLM Harmfulness Perception
@@
The prediction accuracy on the test dataset of three probes is shown in Fig. 2(a). Taking the Llama2-7b model as an example, after the middle layers (around the $10_{th}$ layer), the prediction accuracy of all probes approaches 99%, and the linear probe slightly outperforms the other two. These results demonstrate that the model encodes safety concepts within its hidden states, enabling it to effectively distinguish between safe and harmful prompts. Note that we use various probes to confirm that the separation of harmful and safe prompts in representation space reflects model’s encoding of harmfulness, not due to the structure of a particular probe. Consistent findings across different probes eliminate the influence of probe’s structure, reinforcing that representations capture model’s harmfulness perception.
@@
To further validate whether the directions identified by probes represent safety, we project these direction vectors into vocabulary space and examine the top probability words corresponding for both positive and negative direction vectors on Llama2-7b in Tab. 1. Only few decoded words of the linear probe’s contain affirmation (e.g., ‘Yes’ for positive direction) or refusal (e.g., ‘not’ for negative direction) meaning while other words lack clear interpretability. However, the PCA and cluster probes provide more interpretable results, i.e., the positive direction consistently decodes into affirmation words (e.g., ‘Sure’ and ‘certain’), while the negative direction produces refusal words (e.g., ‘cannot’ and ‘Sorry’). This strongly validates that the directions identified by these probes correspond to safety-related concepts, demonstrating that internal representations contain an inherent sense of safety, which can be effectively captured by probes.
Observation 1. The probing results reveal that aligned LLMs’ representations contain a discernible understanding of safety concept, with specific directions corresponding to affirmation or refusal, indicating an embedded capacity to differentiate between safe and harmful prompts.
Table 1. Direction vectors decoding in vocabulary space, where positive directions often produce affirmation-related words, and negative directions produce refusal-related ones.
@@
5.2.2. Probing Deceptive Harmfulness in Jailbreak Prompts
@@
The probability on predicting as safe by the well-trained safety probes on jailbreak representations from each layer are shown in Fig. 2(b). We observe that most jailbreak methods maintain a high probability of being predicted as safe by the probes in the few early layers (i.e., around the first 10 layers). This suggests that although jailbreak strategies vary, their core mechanism lies in deceiving model’s perception of harmfulness, i.e., making the model treat jailbreak prompts containing malicious intent as safe.
@@
However, different jailbreak methods vary in their ability to deceive model’s harmfulness detection in deeper layers (i.e., around the last 10 layers). For example, in Llama2-7b (the first row in Fig. 2(b)), rule-based and demonstration-based methods such as ReNellm, CodeChameleon, and DeepInception are more deceptive than other methods, identified as safe across all layers. In contrast, evolution-based and multi-agent methods like Autodan and PAIR are less deceptive. Although these methods initially deceive the model into perceiving their prompts as safe in early layers, they are still recognized as harmful till the final layers.
@@
Observation 2. Jailbreak prompts bypass safeguard by deceiving model’s harmfulness perception but varying in effectiveness. Rule-based and demonstration-based methods are more deceptive, maintaining safe predictions even in the final layers, whereas other methods gradually identify their harmful intent as the model processes deeper layers.
@@
Figure 3. Decoding the jailbreak representation of each layer in Llama2-7b in vocabulary space.
5.2.3. Decoding the Jailbreak Representations
We first investigate how Llama2-7b generates tones when processing normal safe and harmful prompts by projecting the corresponding representations from each model layer into the vocabulary space. Since most words decoded from representations in the early few layers are meaningless, we only display and analyze the decoding results of later layers, illustrated in Fig. 3. Rrepresentation of safe prompts in the middle layer (around the $18_{th}$ layer) consistently maps to affirmative words such as ‘Sure’ with high probability, without any refusal words like ‘Sorry’ appearing throughout the process. For harmful prompts, the trend is reversed, with the decoded words consistently carrying meanings of refusal.
@@ 5.2.3. Decoding the Jailbreak Representations
Then we explore how the tone generation for jailbreak prompt responses differs from normal prompts. For rule-based and demonstration-based methods, such as the DeepInception and ReNellm prompt in Fig. 3, the initial word predictions for most prompts tend to produce words that align with the instruction, including task-specific terms like ‘def’ and ‘Layer’, as well as affirmative words like ‘Here’ and ‘Sure’. Notably, no refusal words appear throughout the decoding process. The decoding behavior of these jailbreak prompts closely resembles that of safe prompts, indicating that these prompts successfully bypass model’s harmfulness perception at representation level.
@@
For some prompts of gradient-based jailbreak methods, the generation tone is not consistent. The example GCG prompt in Fig. 3 whose decoding results match this pattern, where affirmation words (such as ‘Sure’) and refusal words (such as ‘Sorry’) both appear simultaneously in the top-5 predictions, with affirmation words dominating with significantly higher probabilities in the last few layers. This suggests that while these jailbreak prompts do not fully deceive the model, i.e., refusal words still appear with high probability, model’s representations are sufficiently perturbed, with affirmation words outweighing refusal words in the final layers.
@@
In contrast, for some rule-based and multi-agent-based jailbreak prompts (e.g., PAIR in Fig. 3), no affirmation words are decoded from hidden states, while refusal words maintain high probability from the middle layers to the final layer, closely resembling the behavior of harmful prompts. This indicates that these jailbreak prompts fail to deceive the model at the first token generation, i.e., the model recognizes the harmful nature of these prompts.
@@
Observation 3. Highly deceptive prompt representations maintain consistent tone with predominantly affirmative highest-probability decoded tokens, while less deceptive prompts lack such consistency, containing both affirmative and negative tokens.
@@
Figure 4. Average refusal score for each attention head in Llama2-7b and Vicuna1.5-7b when responding to harmful and safe prompts, where L21H14 contributes most to refusal, and L26H04 contributes most to affirmations.
5.3. Results of Circuit Analysis
In this section, we present the experimental results and findings on the jailbreak mechanism at the circuit level. Specifically, we conduct experiments from the following perspectives: ➀ Key Components Location, ➁ Effect of Jailbreak Prompts on Key Activations, and ➂ Impact of Model Scale and Fine-tuning.
For ➀ ey components location, we firstly analyze the behavior of important components in transformer based LLM (i.e., all attention heads and MLP layers) on harmful and safe prompts and identify the components that have significant contribution to model safeguard via logit difference attribution introduced in Alg. 1. Then we convert the activation of key components into vocabulary space to validate whether the functionality of identified components is associated with generation safety. For ➁ effect of jailbreak prompts on key activations, we investigate how these key activations are impacted by the jailbreak prompts. Finally, for ➂ impact of model scale and fine-tuning, we compare the model behaviors towards jailbreak prompts in terms of representation and circuit between Llama-2 models and Vicuna-1.5 models. For visual clarity, only the results on the Llama2-7b and Vicuna1.5-7b models are presented in this section while more detailed experimental results on other models are in Appendix. B.
5.3.1. Key Components Location
To identify the key components contributing to refusal responses, we first extract words with the highest predicted probabilities along the safety direction vectors from the probes at final layer for both safe and harmful prompts. For Llama2-7b, the words are ‘Sure’ and ‘Sorry’, while for Vicuna1.5-7b, they are ‘Title’ and ‘Unfortunately’. Then we visualize the direct logit attribution defined in Eq. (3) of each attention head and MLP layer in Llama2-7b and Vicuna1.5-7b, where the results are shown in Fig. 4 and Fig. 5(a), respectively.
As shown in Fig. 4(a), only a small number of attention heads have a significant impact on safety in Llama2-7b. Specifically, the $14_{th}$ attention head in the $21_{st}$ layer (L21H14) in Llama2-7b has a strong contribution to refusing harmful instructions which we denote as the refusal signal head (i.e., $S_{-}$). While the $4_{th}$ head in the $26_{th}$ layer (L26H04) appears to enhance the affirmation response, we refer to it as the affirmation signal head (i.e., $S_{+}$). Moreover, as shown in Fig. 5(a), the $22_{nd}$ MLP in Llama2-7b plays the most critical role in producing affirmative responses (i.e., the reverse of refusal), and the contribution of individual MLP layers to the refusal score on harmful prompts is relatively minor. Early studies (Michel et al., 2019; Voita et al., 2019) on the importance of transformer components highlighted the model’s sparsity, meaning that removing a substantial fraction of attention heads may not degrade model performance. Our findings reinforce the previous observations, confirming that only a small subset of model components are crucial for generation safety.